OpenStudy (anonymous):

limits x approaches 5 ((sqrt x + sqrt 5)/(x-5))

OpenStudy (anonymous):

\[\lim_{x \rightarrow 5} \frac{ (\sqrt{x}-\sqrt{5) } }{ x-5 }\]

OpenStudy (goformit100):

Can I help you Ma'am ?

OpenStudy (anonymous):

is the answer to this \[\frac{ 1 }{\sqrt{10}}\]

OpenStudy (goformit100):

Show your working

OpenStudy (anonymous):

multiply both top and bottom by reciprocal (Sqrt x + Sqrt 5) then your left with x-5 over (x-5)(Sqrt x+Sqrt 5). the x-5 cancel and then its 1/(Sqrt x+sqrt5). then you take the lim and its\[\frac{ 1 }{ \sqrt{5}+\sqrt{5} } = \frac{ 1 }{ \sqrt{10} }\] right?

OpenStudy (goformit100):

Yes it seems to me correct, great work. Keep it up!

OpenStudy (anonymous):

Thank you :)

OpenStudy (goformit100):

My Pleasure. Ask help when ever you like to, I am there to help you

OpenStudy (anonymous):

Ok what about \[\lim_{x \rightarrow 0}\frac{ sinx }{ x ^{2}+2x}\] do i just do the same thing by multiplying by reciprocal?

OpenStudy (goformit100):

it's called commponendo and divinendo.

OpenStudy (anonymous):

what is?

OpenStudy (goformit100):

multiplying by reciprocal

OpenStudy (anonymous):

ok.. how would i start this?

OpenStudy (goformit100):

Ok show your first step

OpenStudy (anonymous):

i would factor the bottom\[\frac{ \sin x }{ x(x+2) }\] and \[\frac{ \sin x }{ x }= 1\] so then the answer is \[\frac{ 1 }{ 2 }\]

OpenStudy (goformit100):

your answer seems correct to me.

OpenStudy (anonymous):

Thank you soo much

OpenStudy (goformit100):

My Pleasure again